# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ Appendix C Mean-field analyses of the low-temperature Higgs phases
The kinetic term (100) is minimized by the condition $\Phi_{\bm{x}}=U_{{\bm{x}},\mu}\Phi_{{\bm{x}}+\hat{\mu}}$, which implies $V(\Phi_{\bm{x}})=V(\Phi_{{\bm{x}}+\hat{\mu}})$. It follows that minimum configurations are obtained by minimizing the potential energy on each lattice site. Therefore, below we neglect the ${\bm{x}}$ dependence of the field. It is trivial to verify that the symmetry properties of the minimum potential configurations do not depend on $r$ and $u$ and therefore, we only report explicit results for the unit-length limit in which ${\rm Tr}\,\Phi^{\dagger}\Phi^{\phantom{\dagger}}=1$.
@@
Here $\phi$ is a complex phase and $F^{af}$ is a unitary $matrix$. If we substitute these expressions in the relation $\Phi_{\bm{x}}=\Pi_{{\bm{x}},\mu\nu}\Phi_{\bm{x}}$ that follows from the minimization of the kinetic energy, one can verify that the plaquette $\Pi_{{\bm{x}},\mu\nu}$ has $q$ unit eigenvalues. Thus, for $N_{f}\geq N_{c}$, $\Pi_{{\bm{x}},\mu\nu}=1$ and the gauge variables are gauge equivalent to the trivial configuration, i.e., $U_{{\bm{x}},\mu}=V_{\bm{x}}^{\dagger}V_{{\bm{x}}+\hat{\mu}}$ where $V_{\bm{x}}\in$ SU($N_{c}$). These results allow us to exclude the presence of topological transitions, related to the residual gauge symmetry in the Higgs phase, see Sec. 3.5.3. Indeed, for $N_{f}\geq N_{c}$ there is no residual gauge symmetry, while for $N_{f}<N_{c}$, gauge modes are controlled by a residual SU($N_{c}-N_{f}$) gauge theory that does not undergo finite-temperature transitions.
@@
where $Y=I_{N_{c}}\oplus 0$ is an $N_{f}\times N_{f}$ diagonal matrix in which the first $N_{c}$ elements are one and the other $N_{f}-N_{c}$ elements are zero, and $\widehat{V}=V\oplus I_{N_{f}-N_{c}}$. This Hamiltonian is invariant under the global transformations $F_{\bm{x}}\to F_{\bm{x}}M$, with $M\in\mathrm{U}(N_{f})$, and under the local transformations $F_{\bm{x}}\to W_{\bm{x}}F_{\bm{x}}$ and $\widehat{V}_{\bm{x}}\to\widehat{V}_{\bm{x}}G_{\bm{x}}$, with $W_{\bm{x}}=W^{(1)}_{\bm{x}}\oplus W^{(2)}_{\bm{x}}$ and $G_{\bm{x}}=W^{(1)}_{\bm{x}}\oplus I_{N_{f}-N_{c}}$, where $W^{(1)}_{\bm{x}}\in\mathrm{SU}(N_{c})$, $W^{(2)}_{\bm{x}}\in\mathrm{U}(N_{f}-N_{c})$ ($F_{\bm{x}}$ is unitary so that $F^{{\dagger}}_{\bm{x}}F_{\bm{x}}=I_{N_{f}}$). The global symmetry of the effective model that describes the critical fluctuations is therefore $\mathrm{SU}(N_{f})/[\mathrm{SU}(N_{c})\otimes\mathrm{SU}(N_{f}-N_{c})]$, which corresponds to the global symmetry-breaking pattern ${\rm U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$, where we have also included the unbroken U(1) symmetry.

### v3

@@ Appendix C Mean-field analysis of the low-temperature Higgs phases
The kinetic term (103) is minimized by the condition $\Phi_{\bm{x}}=U_{{\bm{x}},\mu}\Phi_{{\bm{x}}+\hat{\mu}}$, which implies $V(\Phi_{\bm{x}})=V(\Phi_{{\bm{x}}+\hat{\mu}})$. It follows that minimum configurations are obtained by minimizing the potential energy on each lattice site. Therefore, below we neglect the ${\bm{x}}$ dependence of the field. One can easily verify that the symmetry properties of the minimum-potential configurations do not depend on $r$ and $u$, and therefore we only report explicit results for the unit-length limit in which ${\rm Tr}\,\Phi^{\dagger}\Phi^{\phantom{\dagger}}=1$.
@@
If we substitute these expressions in the relation $\Phi_{\bm{x}}=\Pi_{{\bm{x}},\mu\nu}\Phi_{\bm{x}}$ that follows from the minimization of the kinetic energy, one can verify that the plaquette $\Pi_{{\bm{x}},\mu\nu}$ has $q$ unit eigenvalues. Thus, for $N_{f}\geq N_{c}$, $\Pi_{{\bm{x}},\mu\nu}=1$ and the gauge variables are gauge equivalent to the trivial configuration, i.e., $U_{{\bm{x}},\mu}=V_{\bm{x}}^{\dagger}V_{{\bm{x}}+\hat{\mu}}$ where $V_{\bm{x}}\in$ SU($N_{c}$). These results allow us to exclude the presence of topological transitions, related to the residual gauge symmetry in the Higgs phase, see Sec. 3.5.3. Indeed, for $N_{f}\geq N_{c}$ there is no residual gauge symmetry, while for $N_{f}<N_{c}$, gauge modes are controlled by a residual SU($N_{c}-N_{f}$) gauge theory that does not undergo finite-temperature transitions.
@@
where $Y=I_{N_{c}}\oplus 0$ is an $N_{f}\times N_{f}$ diagonal matrix in which the first $N_{c}$ elements are one and the other $N_{f}-N_{c}$ elements are zero, and $\widehat{V}=V\oplus I_{N_{f}-N_{c}}$. This Hamiltonian is invariant under the global transformations $F_{\bm{x}}\to F_{\bm{x}}M$, with $M\in\mathrm{U}(N_{f})$, and under the local transformations $F_{\bm{x}}\to W_{\bm{x}}F_{\bm{x}}$ and $\widehat{V}_{\bm{x}}\to\widehat{V}_{\bm{x}}G_{\bm{x}}$, with $W_{\bm{x}}=W^{(1)}_{\bm{x}}\oplus W^{(2)}_{\bm{x}}$ and $G_{\bm{x}}=W^{(1)}_{\bm{x}}\oplus I_{N_{f}-N_{c}}$, where $W^{(1)}_{\bm{x}}\in\mathrm{SU}(N_{c})$, $W^{(2)}_{\bm{x}}\in\mathrm{U}(N_{f}-N_{c})$ ($F_{\bm{x}}$ is unitary so that $F^{{\dagger}}_{\bm{x}}F_{\bm{x}}=I_{N_{f}}$). Therefore the global symmetry of the effective model that describes the critical fluctuations is $\mathrm{SU}(N_{f})/[\mathrm{SU}(N_{c})\otimes\mathrm{SU}(N_{f}-N_{c})]$, which corresponds to the global symmetry-breaking pattern ${\rm U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$, where we have also included the unbroken U(1) symmetry.